Grade 11General Physics

A constant force acting on a body of mass 3.0 kg changes its speed from 2.0 m/s to 3.5 m/s in
25 s. The direction of the motion of the body remains unchanged. What is the magnitude and direction of
force?

Nandan Thadukkassery
9 Years agoGrade 11

1 Answer

Hardik Kumar
9 Years ago
Mass(m)=3kgInitial velocity(u)=2m/sFinal velocity(v)=3.5m/sTime taken in this change(t)=25secAcceleration=(v-u)/t ( 3.5 -2 )/25=0.06 m/s^2Force = mass × acceleration 3 × 0.06 = 0.18NDirection of force is along the direction of motion.